1=-20 log k value of k
1
Expert's answer
2018-04-10T04:31:12-0400
Let the base of logarithm be C. Then it can be rewritten as follows:
log C = -20 log k, log C = log k^(-20).
Then it can be overwritten like C=k^(-20) and k=C^(-1/20).